can you solve this equation?
y+2(y-5)=2y+2
1y+2y-5=2y+2
3y-5=2y+2
3y=2y+3
y=3

Answer by sdmmadam@yahoo.com(530)

y+2(y-5)=2y+2 ----(1)
1y+2y-10=2y+2
3y-10=2y+2
(3y -2y) = 2+10
y = 12
Note: 2 multiplied by (-5) is -10 and not (-5)
Verification: y = 12 in (1)
LHS = y+2(y-5)
= 12+2X(12-5)
=12+2X7
=12+14
=26
RHS = 2y+2
=2X12+2
=24+2
=26=LHS